Project

General

Profile

NDOS Commissioning Meeting 2011/06/23

Attendance: Jon, Athans, Andrew, Mark Mathis, Martin, Susan, Jan, Dominick, Gary, Jarek, Enrique, Mat Muether, Xinchun,

Agenda:

  1. Status of NuMI Beam (Jon)
  2. Status of "Cool Running" (Athans, Ron, Peter)
  3. Status of APD visual inspection (Mat, Denis)
  4. DataCheck Update (UMN)
  5. DAQ/Electronics Update & Plans (Leon, Peter)
  6. DCS Plans (Athans)
  7. AOB

Status of NuMI Beam

  • NuMI has resumed LE running. No news re: the leak in the target.

Status of "Cool Running"

  • Athans:
    • Ron is working on maintaining enabled TECCs whenever FEBs are reprogrammed.
    • GUI development is in progress. There will be a meeting tomorrow morning to discuss how to implement the GUI, features, etc.
  • Jon: we attempted to run with 22 cooled APDs on diblock 1 on Tuesday and found that doing a DSO scan resets the FEBs and disables cooling. So were not able to get new pixels masks/thresholds for these APDs. We then decided to start a run with these APDs warm, and then enabled cooling after a few minutes. This seemed to generate a lot of noise to the point where the DCMs were dropping data. Leon later recommended we lower HV on cooled channels as a temporary fix. This helped with all but 4 channels. In the end we disabled cooling and the APDs went back to operating as they did before.
  • Until issue with running DSO scans is solved, we will not be able run with cooled APDs. Ron's work is critical

Status of APD visual inspection

  • Inspection has been completed.
  • Removed 152 APDs that had either obvious cooling problems or installation issues.
  • All have been shipped to CalTech, where they will be tested and baked where necessary.
  • Jon: any part of the detector that had heavier concentration? Mat: 20 APDs were removed from MC. All upstream of diblock 1 had large number of APDs removed.
  • Hardware dB has been updated to reflect these changes to the detector.
  • Most of the problems are traced back to installation errors.
  • Susan: any idea where the oil is coming from on the ones that were marked as "oily"? Mat: not really. Discussion ensued, could it be that oil is coming out of optical connection? The oily APDs seem to all come from the side of the detector.

DataCheck Update

DAQ Update/Plans

  • Peter: has been working on cleaning up OnMon. Found bug in onmon viewer, seems to be more stable now.
  • File transfer will be tested tomorrow morning. Should have no impact on data collection.

DCS Plans

  • Nothing beyond the current work.

AOB

  • Enrique reported that the NDSB door seems to be unlocked, he was able to get into the NDSB without scanning his id.